# LlavaGuard: An Open VLM-based Framework for Safeguarding Vision Datasets and Models

### v2

@@ 5 LlavaGuard Model Suite
To elicit an understanding of safety risks according to a policy, we developed LlavaGuard by leveraging the foundational capabilities of pre-trained VLMs. To evaluate base models for safeguarding, we develop a prompt-response setup. Lastly, we present the training procedure to obtain LlavaGuard.
Prompt-Response Setup. Next to our safety taxonomy, which serves as the default policy prompt (for details, see App. A), a reliably structured output that can be parsed automatically is essential for evaluating visual content at scale. Thus, we task the VLM to assess a given input image against the defined policy by generating a JSON-formatted assessment comprising the following three fields (cf. Fig. 2). First, the (1) safety rating indicates the outcome of the assessment, which can be either Unsafe if the image requires further examination or Safe if it meets the policy standards according to the taxonomy. The (2) category specifies the respective safety category of the taxonomy best describing the image (see Tab. 1). Lastly, the (3) rationale provides a natural language description of the image contents with respect to the policy and selected safety category.
Measuring Policy Responsiveness. To ensure a safeguard’s effectiveness across diverse scenarios, it must flexibly adhere to various policies. We assess this capability using the Policy Exception Rate (PER), which measures the percentage of correctly solved policy exception samples defined through data augmentation.
@@
LlavaGuard Training. We make our training pipeline publicly available. For hyper-parameters and additional details on model tuning, see App. B. We present two different LlavaGuard models with varying model sizes, namely 0.5B and 7B parameters. Both models build upon the Llava-OneVision models (Li et al., 2024b) of their respective sizes.
Inference speed. Speed is a crucial factor when annotating images at scale. The 0.5B model is 347% faster than the 7B model, with an inference time of 0.075s/sample as measured on a single A100 GPU. This speed advantage becomes even more pronounced as GPU memory increases.
@@ 6 Experimental Evaluation
We start with a comprehensive evaluation of LlavaGuard. First, we present qualitative examples to understand potential use cases and the quality of LlavaGuard’s safety assessments. Next, we explore the inferior performance of state-of-the-art safeguards and VLMs compared to LlavaGuard, before turning to one of LlavaGuard’s practical applications in the following section.
Qualitative Results. Fig. 2 presents qualitative examples from the LlavaGuard test set. LlavaGuard’s assessments include safety rating, category, and rationale. Our model not only assigns accurate ratings and categories but also demonstrates transparent policy-following capabilities within the rationales. Specifically, LlavaGuard utilizes the defined policies to assess images, clearly explaining how and why each image complies with or violates the risk guidelines. Furthermore, when the safety policy is modified, LlavaGuard appropriately adjusts its assessments—changing the safety rating from unsafe to safe—and provides solid justifications for these changes in the rationale. In App. Fig. 7, we extend our qualitative evaluation by comparing the assessments of LlavaGuard and corresponding Llava base model using additional images from our test set. We observe that while the base model effectively identifies the content of the images, it fails to adhere to the safety policy. In fact, the policy appears to have no major impact on the base model’s evaluations; it does not account for the policy guidelines within its rationale, nor is it able to adjust its assessment when the policy is modified. In contrast, LlavaGuard provides consistent assessments across these examples and continues to demonstrate strong policy-following capabilities, providing well-grounded reasoning using the risk guidelines of the relevant safety category. Additionally, it demonstrates excellent responsiveness to policy changes.
@@
Table 2: Performance comparison of LlavaGuard with Llava-OV (Li et al., 2024c), GPT-4o, OpenAI Omni-Mod (OpenAI, 2024b), LlamaGuard-3-11B (Chi et al., 2024), and ImageGuard (Li et al., 2025) on the held-out test set. We report balanced accuracy, recall, precision, and policy exception score. LlavaGuard substantially outperforms its baselines and SOTA models. Best values are bold, while runner-up is underlined.
Models Open Acc Recall Preci- PES
(%$\uparrow$ (%$\uparrow$ sion (%$\uparrow$ (%$\uparrow$
VLMs Llava-OV-0.5B 52.00 04.23 90.00 68.07
@@
GPT-4o222experiments performed with gpt-4o-2024-11-20 72.92 55.99 81.05 77.29
@@
LlavaGuard-7B 90.84 91.39 87.97 89.85
Empirical Results. In Tab. 2, we expand upon previous qualitative findings and compare LlavaGuard’s performance on our held-out test set to its baseline VLMs and state-of-the-art safeguards. First, both LlavaGuard models consistently outperform their baselines, improving balanced accuracy by more than 30% compared to Llava-OV (Li et al., 2024c). Furthermore, while multiple other SOTA safeguards demonstrate basic safety understanding in images, their overall ability to evaluate safety is strongly limited. For example, the Meta’s moderation tool Llama-Guard-3-11B-Vision (Chi et al., 2024), has an almost negligible recall, misclassifying nearly all images as ‘safe,’ rendering it unreliable for assessing image safety. Even more strikingly, despite being explicitly designed for this task, both ImageGuard (Li et al., 2025) and OpenAI’s omni moderation (OpenAI, 2024b) achieve only around 70% accuracy–significantly underperforming compared to LlavaGuard. Additionally, LlavaGuard is the only model that demonstrates the desired ability to discern and reject unsafe visual content as evidenced by its recall performance along with high accuracy.
@@ 7 LlavaGuard: Applied Use Cases
Dataset Auditing. In the context of dataset auditing (Gebru et al., 2021), LlavaGuard serves as an annotation tool to identify and document risks associated with unsafe content in large-scale dataset. This helps ensure the integrity and safety of data and downstream AI models (Schramowski et al., 2022; Birhane et al., 2021).
@@
Fig. 4 illustrates that LlavaGuard assigns 105k images out of 1.3M from ImageNet to one of the 9 safety categories. Among these, 20k instances (19% of the subset and 1.5% of the entire ImageNet) violate the safety policy, thus are rated as unsafe. On the other hand, the vast majority of ImageNet (98.5%) adheres to the safety standards and was rated as safe. Many images fall under category O6: Weapons or Substance Abuse which is a result of ImageNet classes ‘assault_rifle’, ‘tank’, and ‘rifle’. Yet, LlavaGuard clearly distinguishes between guideline-violating images (only 11k out of 36k). These findings, are in line with previous works (Schramowski et al., 2022; Birhane & Prabhu, 2021) which have also identified a substantial number of potentially unsafe images in ImageNet. For example, Schramowski et al.’s classifier flagged over 40k unsafe images. However, upon manual inspection, we found their classifier to be more conservative, failing to differentiate between benign depictions of weapons and illegal ones.
@@
In Fig. 4(b), we present examples of unsafe images from ImageNet. These samples are clearly unsafe and violate the safety policy. In addition, the assigned safety categories are well aligned with the depicted content. These examples underscore a critical challenge with large-scale datasets: while the general use of these images may be problematic, the human-assigned labels are often even more questionable. In more detail, the labels assigned often do not align with the core content of the image. For instance, the image at bottom center is labeled as ‘bath tub’, yet it primarily displays explicit nude content (O4). Associations like these can lead to spurious, harmful correlations in models trained on this data. These findings underline the need for advanced auditing tools like LlavaGuard in data curation and preprocessing pipelines, especially when handling data at scale where full manual annotation is not feasible.
Model Safeguarding. While dataset auditing can lead to safer models, implementing adequate safeguards during deployment remains crucial. Consequently, we considered StableDiffusion-v1.5 (SD1.5) (Rombach et al., 2022), a model known for its susceptibility to generating unsafe material (Schramowski et al., 2023). We leverage the distilled inappropriate image prompts (I2P) benchmark (Brack et al., 2023b) to elicit the generation of potentially problematic material and subsequently analyze the generated images with LlavaGuard. These prompts (1.1k in total) are specifically designed to evade classical input filters to result in unsafe images. We generated 10 images for each prompt, resulting in 11k images.
@@
To validate LlavaGuard’s assessments, we manually probed the generated images and largely agree. Thus, confirming observations of previous works (Birhane et al., 2021; Schramowski et al., 2023; Brack et al., 2023b) that sexually explicit and nude imagery of women is remarkably easy to produce with seemingly safe prompts. This behavior urges more research into safe generative models and the development of safety guardrails.
Table 3: Agreement rate of human users with LlavaGuard assessments. The high agreement with human assessments shows that LlavaGuard accurately captures safety in vision content across all three dimensions.
human agreement rating category rationale
88% 87% 81%
Measuring Human Agreement with LlavaGuard. To extend previous evidence, we asked human users to annotate LlavaGuard assessments from a broad array of applied use cases. The assessments of LlavaGuard are sampled across a diverse range of datasets—including both real (CC12M (Changpinyo et al., 2021a), COCO (Lin et al., 2014a), ImageNet (Deng et al., 2009)) and synthetic (I2P-generated images (Brack et al., 2023b) by StableDiffusion 1.5, Stylebreeder (Zheng et al., 2024)) datasets—, providing a strong coverage across domains. For further details, see App. J. As shown in Tab. 3, LlavaGuard demonstrates a high level of agreement with human assessments, particularly in safety ratings and category classifications. The agreement is naturally slightly lower for the more complex rationales. These results underscore LlavaGuard’s effectiveness in delivering high-quality, human-aligned evaluations across various applications.
@@ 8 Conclusion
We introduced a novel framework for vision safeguard models, including a safety risk taxonomy for assessing the safety of images alongside a human-annotated safety dataset labeled based on this taxonomy. LlavaGuard goes beyond rigid classifications and provides assessments that include violated categories and detailed rationales. Our empirical results show that LlavaGuard serves as a strong cornerstone for VLM-based safeguarding vision datasets and models. For future work, LlavaGuard would generally benefit from extending its training and test data, specifically with synthetic content. Another promising area for exploration involves extending the categories to encompass bias assessment to promote fairness.
Limitations. During the tuning process of LlavaGuard models, human supervision was applied solely to category and rating entries, while the rationales were generated synthetically. Additionally, the annotation of our dataset was largely guided by the default policy outlined in App. A. While we incorporated policy permutations during training to accommodate diverse policy specifications, we encourage future work to explore annotations that consider varying policies. The tradeoff between computational cost and performance is an important consideration, especially when auditing large-scale datasets and run-time monitoring generative models. To address this, we provide both smaller (0.5B) and larger (7B) checkpoints to accommodate varying requirements. We recognize that this work’s safety taxonomy provides foundational coverage of the safety categories, offering opportunities for further expansion and refinement. As previously mentioned, safety is highly context- and situation-dependent, which makes a single, universal definition and taxonomy often impractical. Yet, we argue that, much like in law-making, our taxonomy adopts a reasonable approach by defining a set of general safety rules. Moreover, LlavaGuard already demonstrates remarkable capabilities in handling different policies, as evidenced by its high PES values. Moreover, as more powerful open VLMs become available, it is expected that LlavaGuard will continue to improve, given that it is agnostic to its underlying VLM.
@@
Acknowledgements
We acknowledge support of the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation), the hessian.AISC Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), and the German Research Center for AI (DFKI). Further, this work benefited from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215), the Hessian research priority program LOEWE within the project WhiteBox, the HMWK cluster projects “Adaptive Mind” and “Third Wave of AI”, and from the NHR4CES.
@@ Acknowledgements
We blurred faces in presented images of the main text to protect privacy.
@@ Appendix C LlavaGuard: Further Dataset Audits
Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafeO2O2 Refer to captionSafeUnsafeNAO1
Refer to captionSafeUnsafe05O7 Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafeNAO1
Refer to captionUnsafeO1SafeO1 Refer to captionUnsafeO8SafeO1 Refer to captionUnsafeO2UnsafeO2
Refer to captionUnsafeO3SafeO3 Refer to captionUnsafeO4SafeO3 Refer to captionUnsafeO2UnsafeO2O1
(c) Illustrative examples from ImageNet depicting safe (left) and unsafe (right) images. Safety evaluations (category and rating) from LlavaGuard (blue) and Llava (orange) are provided in the corners.
Figure 8: LlavaGuard vs. Llava in-the-wild: We provide a quantitative (a and b) and qualitative (c) comparison, auditing ImageNet (1.3M images) with our default taxonomy. (a) and (b) report quantitative results encompassing overall category detections and the portion classified as unsafe, for LlavaGuard and Llava, respectively. Overall, Llava (b) flags more images as unsafe and categorizes most of them into O1. LlavaGuard (a), on the other hand, is able to find unsafe images across all categories. When further examining images in (c), one can observe a stark differences in safety annotations. (left) depicts safe and (right) unsafe examples from ImageNet, with safety evaluations (category and rating) from LlavaGuard (blue) and Llava (orange) in the corners. LlavaGuard correctly assigns safety ratings and categories, while Llava’s evaluations are inaccurate. Its safety categories are flawed (O1 is overused), and the safety ratings are incorrect according to the policy. Generally, Llava flags more images as unsafe, but is only able to detect two out of six unsafe images (c) and misclassifies multiple safe images as unsafe. This suggest a superior performance of LlavaGuard and the limitations of baseline Llava.

### v3

@@ 5 LlavaGuard Model Suite
To elicit an understanding of safety risks according to a policy, we developed LlavaGuard by leveraging the foundational capabilities of pre-trained VLMs. To evaluate base models for safeguarding, we develop a prompt-response setup. Lastly, we show how to train LlavaGuard.
5.1 Prompt-Response Setup
Next to our safety taxonomy, which serves as the default policy prompt (for details, see App. A), a reliably structured output that can be parsed automatically is essential for evaluating visual content at scale. Thus, we task the VLM to assess a given input image against the defined policy by generating a JSON-formatted assessment comprising the following three fields (cf. Fig. 2). First, the (1) safety rating indicates the outcome of the assessment, which can be either Unsafe if the image requires further examination or Safe if it meets the policy standards according to the taxonomy. The (2) category specifies the respective safety category of the taxonomy best describing the image (see Tab. 1). Lastly, the (3) rationale provides a natural language description of the image contents with respect to the policy and selected safety category.
5.2 Policy Responsiveness
To ensure a safeguard’s effectiveness across diverse scenarios, it must flexibly adhere to various policies. We assess this capability using the Policy Exception Rate (PER), which measures the percentage of correctly solved policy exception samples defined through data augmentation.
@@ 5.2 Policy Responsiveness
5.3 LlavaGuard Training
Detailed descriptions of the employed hyperparameters and model tuning procedures are provided in App. B. In our experiments, we introduce two versions of LlavaGuard with model sizes of 0.5B and 7B parameters, both of which are based on the corresponding Llava-OneVision architectures(Li et al., 2024b). In addition, we present two QwenGuard variants—comprising 3B and 7B parameters—built upon the Qwen2.5-VL models (Wang et al., 2024; Bai et al., 2023) of matching scale.
Inference speed.
Speed is a crucial factor when annotating images at scale. The 0.5B model is 347% faster than the 7B model, with an inference time of 0.075s/sample as measured on a single A100 GPU. This speed advantage becomes even more pronounced as GPU memory increases.
@@ 6 Experimental Evaluation
We begin with a comprehensive evaluation of LlavaGuard. First, we present qualitative examples to illustrate potential use cases and to assess the quality of LlavaGuard’s safety evaluations. Next, we analyze the limitations and inferior performance of state-of-the-art (SOTA) safeguards and VLMs compared to LlavaGuard. Finally, we demonstrate one of LlavaGuard’s practical applications in the subsequent section, highlighting its effectiveness and versatility in real-world scenarios.
6.1 Qualitative Results
Fig. 2 presents qualitative examples from the LlavaGuard test set. LlavaGuard’s assessments include safety rating, category, and rationale. Our model not only assigns accurate ratings and categories but also demonstrates transparent policy-following capabilities within the rationales. Specifically, LlavaGuard utilizes the defined policies to assess images, clearly explaining how and why each image complies with or violates the risk guidelines. Furthermore, when the safety policy is modified, LlavaGuard appropriately adjusts its assessments—changing the safety rating from unsafe to safe—and provides solid justifications for these changes in the rationale. In App. Fig. 7, we extend our qualitative evaluation by comparing the assessments of LlavaGuard and corresponding Llava base model using additional images from our test set. We observe that while the base model effectively identifies the content of the images, it fails to adhere to the safety policy. In fact, the policy appears to have no major impact on the base model’s evaluations; it does not account for the policy guidelines within its rationale, nor is it able to adjust its assessment when the policy is modified. In contrast, LlavaGuard provides consistent assessments across these examples and continues to demonstrate strong policy-following capabilities, providing well-grounded reasoning using the risk guidelines of the relevant safety category. Additionally, it demonstrates excellent responsiveness to policy changes.
@@ 6.1 Qualitative Results
Table 3: Performance comparison of LlavaGuard and alternative vision-safeguards (including VLM baselines and SOTA moderation tools) on the held-out test set. We report balanced accuracy, recall, precision, and policy exception score (PES). LlavaGuard substantially outperforms both open-source and proprietary baselines. Best values are bold, while runner-up is underlined; higher is better; in [%].
Models Open Accuracy Recall Precision PES
General-Purpose VLMs Llava-OV-0.5B 52.00 04.23 90.00 68.07
@@
InternVL2.5-1B 50.60 88.06 44.03 11.78
InternVL2.5-8B 61.27 31.28 73.68 65.34
InternVL2.5-78B 66.92 46.11 74.44 66.79
Qwen2.5-VL-3B 68.09 79.72 58.69 30.92
Qwen2.5-VL-7B 67.58 49.17 73.14 63.56
Qwen2.5-VL-72B 70.84 60.00 71.76 60.12
QVQ-72B-Preview 62.01 25.54 93.42 74.79
GPT-4o111experiments performed with gpt-4o-2024-11-20 72.92 55.99 81.05 77.29
@@
Ours LlavaGuard-7B 90.84 91.39 87.97 89.85
6.2 Empirical Results
In Tab. 3, we expand upon previous qualitative findings and compare LlavaGuard’s performance on our held-out test set to its baseline VLMs and state-of-the-art safeguards. As an additional, lightweight baseline, we report results of a finetuned Siglip2-large (Tschannen et al., 2025) model trained on our dataset (Siglip2Guard). First, both LlavaGuard models consistently outperform their baselines, improving balanced accuracy by more than 30% compared to Llava-OV (Li et al., 2024c). Furthermore, while multiple other SOTA safeguards demonstrate basic safety understanding in images, their overall ability to evaluate safety is strongly limited. For example, Meta’s moderation tool Llama-Guard-3-11B-Vision (Chi et al., 2024), has an almost negligible recall, misclassifying nearly all images as ‘safe,’ rendering it unreliable for assessing image safety. Even more strikingly, despite being explicitly designed for this task, both ImageGuard (Li et al., 2025) and OpenAI’s omni moderation (OpenAI, 2024b) achieve only around 70% accuracy–significantly underperforming compared to LlavaGuard. Additionally, LlavaGuard is the only model that demonstrates the desired ability to discern and reject unsafe visual content as evidenced by its recall performance along with high accuracy.
@@ 7 LlavaGuard: Applied Use Cases
7.1 Dataset Auditing
In the context of dataset auditing (Gebru et al., 2021), LlavaGuard serves as an annotation tool to identify, document, and categorize risks associated with the presence of unsafe and harmful content in large-scale datasets. This helps to ensure the integrity and safety of data and downstream AI models (Schramowski et al., 2022; Birhane et al., 2021).
@@ 7.1 Dataset Auditing
Fig. 4 illustrates that LlavaGuard assigns 105k images out of 1.3M from ImageNet to one of the 9 safety categories. Among these, 20k instances (19% of the subset and 1.5% of the entire ImageNet) violate the safety policy and thus are rated as unsafe. On the other hand, the vast majority of ImageNet (98.5%) adheres to the safety standards and was rated as safe. Many images fall under category O6: Weapons or Substance Abuse which is a result of ImageNet classes ‘assault_rifle’, ‘tank’, and ‘rifle’. Yet, LlavaGuard clearly distinguishes between guideline-violating images (only 11k out of 36k). These findings are in line with previous works (Schramowski et al., 2022; Birhane & Prabhu, 2021), which have also identified a substantial number of potentially unsafe images in ImageNet. For example, Schramowski et al.’s classifier flagged over 40k unsafe images. However, upon manual inspection, we found their classifier to be more conservative, failing to differentiate between benign depictions of weapons and illegal ones.
In Fig. 4(b), we present examples of unsafe images from ImageNet. These samples are clearly unsafe and violate the safety policy. In addition, the assigned safety categories are well-aligned with the depicted content. These examples underscore a critical challenge with large-scale datasets: while the general use of these images may be problematic, the human-assigned labels are often even more questionable. In more detail, the labels assigned often do not align with the core content of the image. For instance, the image at the bottom center is labeled as ‘bath tub’, yet it primarily displays explicit nude content (O4). Associations like these can lead to spurious, harmful correlations in models trained on this data. These findings underline the need for advanced auditing tools like LlavaGuard in data curation and preprocessing pipelines, especially when handling data at scale where the full manual annotation is not feasible.
Downstream Performance on Filtered Dataset.
To evaluate the impact of safety filtering on downstream visual recognition, we trained a ResNet-50 model (for 50 epochs, using AdamW with a learning rate of 0.001) from scratch on both the original ImageNet dataset and a LlavaGuard-filtered version, in which approximately 20,000 images ($\sim$1of the data) were removed. The overall classification performance remained virtually unchanged: top-1 accuracy was $67.2\pm 0.5$ for unfiltered and $67.4\pm 0.4$ for filtered data, while top-5 accuracy was $87.2\pm 0.6$ and $87.1\pm 0.6$, respectively. Notably, LlavaGuard removed up to 55% of samples from certain classes, such as “assault rifle,” “army tank,” “missile,” and “syringe.” Restricting evaluation to the ten most heavily filtered classes, we observed a more pronounced accuracy drop: top-1 accuracy decreased from $53.6\pm 3.5$ to $46.9\pm 4.8$, and top-5 from $82.4\pm 3.1$ to $77.3\pm 3.6$. These results demonstrate that, when applied carefully, safety filtering can be implemented with minimal effect on aggregate downstream performance, although substantial changes in class distribution may still impact specific categories.
@@
7.2 Model Safeguarding
While dataset auditing can lead to safer models, implementing adequate safeguards during deployment remains crucial. Consequently, we considered StableDiffusion-v1.5 (SD1.5) (Rombach et al., 2022), a model known for its susceptibility to generating unsafe material (Schramowski et al., 2023). We leverage the distilled inappropriate image prompts (I2P) benchmark (Brack et al., 2023b) to elicit the generation of potentially problematic material and subsequently analyze the generated images with LlavaGuard. These prompts (1.1k in total) are specifically designed to evade classical input filters to result in unsafe images. We generated 10 images for each prompt, resulting in 11k images.
@@ 7.2 Model Safeguarding
To validate LlavaGuard’s assessments, we manually probed the generated images and largely agreed. Thus, confirming observations of previous works (Birhane et al., 2021; Schramowski et al., 2023; Brack et al., 2023b) that sexually explicit and nude imagery of women is remarkably easy to produce with seemingly safe prompts. This behavior urges more research into safe generative models and the development of safety guardrails.
7.3 Measuring Human Agreement with LlavaGuard
To extend previous evidence, we asked human users to annotate LlavaGuard assessments from a broad array of applied use cases. The assessments of LlavaGuard are sampled across a diverse range of datasets—including both real (CC12M (Changpinyo et al., 2021a), COCO (Lin et al., 2014a), ImageNet (Deng et al., 2009)) and synthetic (I2P-generated images (Brack et al., 2023b) by StableDiffusion 1.5, Stylebreeder (Zheng et al., 2024)) datasets—, providing a strong coverage across domains. For further details and results, see App. J. Annotators reported a high level of agreement with LlavaGuard, particularly in safety ratings and category classifications. Specifically, we observed agreement for 88% of ratings, 87% of category assignments, and 81% of generated rationales, respectively. The agreement is naturally slightly lower for the more complex rationales. These results underscore LlavaGuard’s effectiveness in delivering high-quality, human-aligned evaluations across various applications.
@@ 8 Conclusion
We introduced a novel framework for vision safeguards, including a safety risk taxonomy for assessing the safety of images alongside a human-annotated safety dataset labeled based on this taxonomy. LlavaGuard goes beyond rigid classifications and provides assessments that include violated categories and detailed rationales. Our empirical results show that LlavaGuard serves as a strong cornerstone for VLM-based safeguarding vision datasets and models. For future work, LlavaGuard would generally benefit from extending its training and test data, specifically with synthetic content. Another promising area for exploration involves extending the categories to encompass bias assessment to promote fairness.
Limitations.
During the tuning process of LlavaGuard models, human supervision was applied solely to category and rating entries, while the rationales were generated synthetically. Additionally, the annotation of our dataset was largely guided by the default policy outlined in App. A. While we incorporated policy permutations during training to accommodate diverse policy specifications, we encourage future work to explore annotations that consider varying policies. The tradeoff between computational cost and performance is an important consideration, especially when auditing large-scale datasets and runtime monitoring generative models. To address this, we provide both smaller (0.5B) and larger (7B) checkpoints to accommodate varying requirements. We recognize that this work’s safety taxonomy provides foundational coverage of the safety categories, offering opportunities for further expansion and refinement. As previously mentioned, safety is highly context- and situation-dependent, which makes a single, universal definition and taxonomy often impractical. Yet, we argue that, much like in law-making, our taxonomy adopts a reasonable approach by defining a set of general safety rules. Moreover, LlavaGuard already demonstrates remarkable capabilities in handling different policies, as evidenced by its high PES values. Moreover, as more powerful open VLMs become available, it is expected that LlavaGuard will continue to improve, given that it is agnostic to its underlying VLM. Finally, while LlavaGuard shows robust performance overall, there remain a small number of challenging cases—such as images close to decision boundaries or images containing complex embedded text—which may still occasionally result in misclassifications (see App.Sec.K).
Acknowledgements
We acknowledge support of the hessian.AI Innovation Lab (funded by the Hessian Ministry for Digital Strategy and Innovation), the hessian.AISC Service Center (funded by the Federal Ministry of Education and Research, BMBF, grant No 01IS22091), and the Centre for European Research in Trusted AI (CERTAIN). Further, this work benefited from the ICT-48 Network of AI Research Excellence Center “TAILOR” (EU Horizon 2020, GA No 952215), the Hessian research priority program LOEWE within the project WhiteBox, the HMWK cluster projects “Adaptive Mind” and “Third Wave of AI”, and from the NHR4CES.
@@ Acknowledgements
We blurred faces in the presented images of the main text to protect privacy.
@@ Appendix C LlavaGuard: Further Dataset Audits
Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafeO2O2 Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafe05O7 Refer to captionSafeUnsafeNAO1 Refer to captionSafeUnsafeNAO1
Refer to captionUnsafeO1SafeO1 Refer to captionUnsafeO8SafeO1 Refer to captionUnsafeO2UnsafeO2 Refer to captionUnsafeO3SafeO3 Refer to captionUnsafeO4SafeO3 Refer to captionUnsafeO2UnsafeO2O1
(c) Illustrative examples from ImageNet depicting safe (top) and unsafe (bottom) images. Safety evaluations (category and rating) from LlavaGuard (blue) and Llava (orange) are provided in the corners.
Figure 8: LlavaGuard vs. Llava in-the-wild: We provide a quantitative (a and b) and qualitative (c) comparison, auditing ImageNet (1.3M images) with our default taxonomy. (a) and (b) report quantitative results encompassing overall category detections and the portion classified as unsafe, for LlavaGuard and Llava, respectively. Overall, Llava (b) flags more images as unsafe and categorizes most of them into O1. LlavaGuard (a), on the other hand, is able to find unsafe images across all categories. When further examining images in (c), one can observe a stark differences in safety annotations. (top) depicts safe and (bottom) unsafe examples from ImageNet, with safety evaluations (category and rating) from LlavaGuard (blue) and Llava (orange) in the corners. LlavaGuard correctly assigns safety ratings and categories, while Llava’s evaluations are inaccurate. Its safety categories are flawed (O1 is overused), and the safety ratings are incorrect according to the policy. Generally, Llava flags more images as unsafe, but is only able to detect two out of six unsafe images (c) and misclassifies multiple safe images as unsafe. This suggest a superior performance of LlavaGuard and the limitations of baseline Llava.